If a circuit contains three 60 Ω resistors in parallel and has a total voltage of 120 V what is the total resistance of the circuit and the total current?   How many electrons does this charge represent?




1
Expert's answer
2021-06-24T16:14:29-0400
"\frac{1}{R}=3\frac{1}{60}\\R=20\Omega\\I=\frac{120}{20}=6\ A"